OpenStudy (caroline1218):

Write an equation in point-slope form for the line through the given point with the given slope. (5, 2); m = 3

OpenStudy (rose4825):

Alright. Let's first determine the equation for point-slope form. Do you know what it is?

OpenStudy (rose4825):

The point slope form is \[y - y _{1} = m(x - x _{1})\]

OpenStudy (rose4825):

You need one set of coordinate points and the value for slope for point slope form. Your coordinate points are (5, 2) and your slope is 3. \[y _{1} = 2\] and \[x _{1} = 5\] the "m" is the same thing as slope. So since m = 3, your slope will also equal 3 (since they are the same thing). Now you have to plug everything in: \[y - 2 = m(x - 5)\]